OpenStudy (anonymous):

Suppose an electronics retailer advertises a 30% discount on laptop computers. If a laptop you are interested in purchasing cost $1299, what will be the amount of the discount and the final sales price after the discount is applied?

OpenStudy (dan815):

30% means 30/100 so basically 30% off means like for everyone 100 dollars you get 30 dollars off, so you only pay 70

OpenStudy (dan815):

so like for 420 dollars you get 30% off... let see now for each 100 dollars u get 30 off There are 4 hunders in 420 and 2 of 10s as 420 = 4*100 + 2*10 so 4*(100-30) + 2(10-3)=4*70+2*7=280+14=294 notice how i knew it would be 3 dollars off the 10 since its 30 dollars off 100 same ratio off... 30%...

OpenStudy (anonymous):

yeah i noticed

OpenStudy (anonymous):

a)$389.70, $909.30, .B) $395.44, $903.56 ,C) $397.60, $901.40.,d). $399.00, $900.00